# Ocean-E2E: Hybrid Physics-Based and Data-Driven Global Forecasting of Extreme Marine Heatwaves with End-to-End Neural Assimilation

### v2

@@ Appendix C Details of Dynamical Core of Ocean-E2E
$C_{t}=\int_{0}^{t}[-({\bf u}(t)+\varepsilon_{gm}{\bf u}_{\theta}({\bf u}(t),C))\nabla C]\mathbf{M}+\dot{\rm S}_{\theta}({\bf A}(t),C)dt+C_{0}$ (28)
Time Integration. As mentioned in Section 3.2, we use second-order Runge-Kutta method to approximate the time integration:
$\tiny C_{(i+1)\Delta t}^{\star}$ $=C_{i\Delta t}+\left[-\left({\bf u}_{i\Delta t}+{\bf u}_{\theta}({\bf u}_{i\Delta t},C_{i\Delta t})\right)\nabla C_{i\Delta t}+\dot{\rm S}_{\theta}(A_{i\Delta t:(i+1)\Delta t},C_{i\Delta t})\right]\Delta t$ (29)
$C_{(i+1)\Delta t}$ $=C_{i\Delta t}+\frac{1}{2}\Bigg{\{}\left[-\left({\bf u}_{i\Delta t}+{\bf u}_{\theta}({\bf u}_{i\Delta t},C_{i\Delta t})\right)\nabla C_{i\Delta t}+\dot{\rm S}_{\theta}(A_{i\Delta t:(i+1)\Delta t},C_{i\Delta t})\right]\Delta t$
$\quad+\left[-\left({\bf u}_{(i+1)\Delta t}+{\bf u}_{\theta}({\bf u}_{(i+1)\Delta t},C_{(i+1)\Delta t}^{\star})\right)\nabla C_{(i+1)\Delta t}^{\star}+\dot{\rm S}_{\theta}(A_{(i+1)\Delta t:(i+2)\Delta t},C_{(i+1)\Delta t}^{\star})\right]\Bigg{\}}$ (30)
In this study, we set $\Delta t=24$ hours and the whole integration time is $t=4$ days.
Appendix D Details of Our Neural Networks
D.1 Multi-scale Convolution
Inspired by [60], this section details the architectures of the neural network modules $\mathbf{u}_{\theta}$ (for subgrid advection) and $\dot{S}_{\theta}$ (for mixing and external source/sink terms) within the Ocean-E2E framework. As shown in Figure 6, both modules leverage the GABlock (Group Attention Block) as their core building unit, designed to efficiently capture multi-scale spatio-temporal features.
@@ D.1 Multi-scale Convolution
The module for approximating the bolus velocity, $\mathbf{u}_{\theta}$, adopts a U-Net-like encoder-decoder architecture. It takes the concatenated ocean surface current $\mathbf{u}\in\mathbb{R}^{B\times 2\times H\times W}$ and SSTA $\mathbf{C}\in\mathbb{R}^{B\times 1\times H\times W}$ (forming $\mathbf{X}_{u}\in\mathbb{R}^{B\times 3\times H\times W}$) as input. The encoder path, composed of $L_{e}$ stages, utilizes downsampling convolutions (e.g., ConvSC) and GABlocks to extract multi-scale features, denoted as $\mathcal{H}_{l}$ for the $l$-th stage. Deeper encoder layers, particularly the bottleneck which processes $\mathcal{H}_{L_{e}}$ with stacked GABlocks, are crucial for capturing non-local spatial dependencies inherent in subgrid effects. The decoder path, with $L_{d}$ stages, symmetrically reconstructs spatial resolution using upsampling convolutions and GABlocks, integrating features from corresponding encoder stages $\mathcal{H}_{L_{e}-l+1}$ via skip connections. A final convolutional layer then maps the decoder’s output to the bolus velocity field $\mathbf{u}_{\theta}\in\mathbb{R}^{B\times 2\times H\times W}$.
The module for parameterizing mixing and external source/sink terms, $\dot{S}_{\theta}$, primarily consists of a stack of $N_{s}$ GABlocks. Its input is the concatenation of atmospheric forcing variables $\mathbf{A}\in\mathbb{R}^{B\times C_{A}\times H\times W}$ and SSTA $\mathbf{C}\in\mathbb{R}^{B\times 1\times H\times W}$ (forming $\mathbf{X}_{s}\in\mathbb{R}^{B\times(C_{A}+1)\times H\times W}$). An initial convolution, InitialConv, first adjusts $\mathbf{X}_{s}$ to a hidden channel dimension $C_{\text{hid}}$. This is followed by $N_{s}$ serially stacked GABlockl ($l=1,\dots,N_{s}$). Let $\mathcal{H}_{l}$ denote the output of the $l$-th GABlock. Finally, an OutputConv layer maps the features from the last GABlock, $\mathcal{H}_{N_{s}}$, to a single-channel output representing the SSTA tendency $\dot{S}_{\theta}\in\mathbb{R}^{B\times 1\times H\times W}$.
These architectures, centered around the versatile GABlock, enable $\mathbf{u}_{\theta}$ and $\dot{S}_{\theta}$ to effectively learn and represent the complex physical processes governing SSTA evolution directly from observational and reanalysis data.
D.2 Neural Data Assimilation Network
Inspired by  [38], we apply an edge-reparameterization and attention-guided network to project the observation to analysis data, which achieves efficent assimilation for real-time application. It primarily consists of four parts, i.e., a convolutional layer (ConvL), a Kirsch-guided reparameterization module (KRM), a channel attention module (CAM), and a spatial attention module (SAM).
D.2.1 Convolutional Layer
@@ D.2.2 Attention Mechanism
[…]
@@
[…]
D.2.3 Kirsch-guided Reparameterization Module
To facilitate efficient assimilation in real-world MHWs forecasting, we employ the Kirsch-guided reparameterization module (KRM) with shared parameters, as proposed by Liu et al. [38]. The KRM is composed of three principal elements: (1) a standard convolutional layer that captures local spatial features; (2) an expand–squeeze convolutional layer that adaptively adjusts the receptive field; and (3) a family of eight directional edge-detection operators that guide the learning and inference of the convolutional kernels. Specifically, the standard convolutional and expand–squeeze operations can be formulated as follows:
@@ D.2.3 Kirsch-guided Reparameterization Module
where $\operatorname{perm}(\cdot)$ denotes the permute operation which exchanges the $1$st and $2$nd dimensions of a tensor, $\operatorname{rep}(\cdot)$ denotes the spatial broadcasting operation. And the predefined eight-direction Kirsch edge filters $K_{i}$ are incorporated into the reparameterization module. To memorize the edge features, the input feature $x_{in}^{k}$ will first be processed by $C\times C\times 1\times 1$ convolution and then use a custom Kirsch filter to extract the feature map gradients in eight different directions. Therefore, the edge information in eight directions can be expressed as follows:
@@ D.2.4 The Assimilation Netwrok
Appendix E Experiments Details
E.1 Evaluation Metric
Definition of MHWs. The [25] definition of MHW categories is applied to detect MHWs from original SST field. More precisely, any interval in which the SST anomalies surpass the 90th percentile for a minimum of five days in a row is classified as an MHW event. Subsequently, the MHWs are divided into three severity categories: moderate, strong, and severe/extreme. These categories are determined by the 90th, 92.5th, and 95th percentiles of the SST, respectively [20]. Note that our results are unaffected by different definitions of the MHW category.
@@ E.1 Evaluation Metric
${\rm RMSE}(t)=\sqrt{\frac{1}{H\times W}\sum_{i,j}\omega_{i,j}(T_{i,j}(t)-\tilde{T}_{i,j}(t))^{2}}$ (58)
Where $T_{i,j}(t)$ and $\tilde{T}_{i,j}(t)$ is the ground truth and prediction of SST anomaly at lead time $t$, $\omega_{i,j}$ is the weighting coefficient and the subscript $(i,j)$ indicates the data at grid point $(i,j)$ while $H$ and $W$ follow the same definition in Section 2. CSI is the metric that evaluate the forecast ability of extreme events and can be expressed as:
@@
${\rm BIAS}(t)=\frac{1}{H\times W}\sum_{i,j}\omega_{i,j}(T_{i,j}(t)-\tilde{T}_{i,j}(t))$ (60)
E.2 Model Training
@@
Appendix F Related Work and Limitations
F.1 Numerical Model based Ocean Forecasting
@@
Traditional MHWs forecasting predominantly relies on numerical models solving oceanic primitive equations. These approaches are typically categorized into two types: (1) seasonal forecasting, which predicts large-scale MHW characteristics including onset timing, duration, and intensity evolution [30, 8], and (2) sub-seasonal forecasting, which resolves finer spatiotemporal variations at the cost of reduced prediction horizons [2, 63]. While physics-based models demonstrate reasonable forecast skill, they face two principal limitations: prohibitive computational costs scaling with model resolution [62, 58], and systematic underestimation of extreme MHW magnitudes [30].
F.2 Deep Learning based Ocean Forecasting
@@
Machine learning (ML) techniques have emerged as promising alternatives for oceanic modeling, offering orders-of-magnitude speedup over conventional methods [37, 22]. Recent efforts have developed spatio-temporal architectures including NMO [59], neXtSIM [16], and OceanVP [50] for regional ocean modeling, alongside specialized models for specific oceanic phenomena such as ENSO [21, 12], MJO [31, 14, 51], and MHWs [49, 30, 52]. Building on advances in atmospheric foundation models [44, 4], the community has recently proposed global ocean foundation models like AI-GOMS [62], Xihe [58], and WenHai [13].
F.3 Deep Learning based Data Assimilation
@@
Deep learning (DL) has shown transformative potential in weather and climate data assimilation (DA), with progress evolving along three directions: (1) Conceptual validation using idealized systems like Lorenz models and shallow water equations [7, 1, 15, 33]; (2) Component replacement within conventional DA frameworks, including physical parameterization [23], observation operators [35, 54], and optimization cost functions [41, 61]; (3) End-to-end DA system replacement, exemplified by FengWu-Adas and DiffDA that implement complete DL-based assimilation pipelines using ERA5 reanalysis data [11, 27].
Appendix G Additional Results
G.1 Additional Comparison with S2S System
@@
In this section, we are going to show additional comparison results with S2S system. To be specific, we take the S2S analysis field as the groundtruth, and use it to initialize our Ocean-E2E. The atmospheric boundary condition is acquired through one of the state-of-the-art AI-driven weather forecast model: OneForecast [17]. As shown in Table 7, our model still performs better than S2S.
@@
G.2 Visualization of Global MHWs Simulation
In this section, we visualize the global MHWs forecast result of our Ocean-E2E and S2S, as shown in Figure 7-10. We plot the global SSTA field, where darkred parts indicate strong MHWs.
G.3 Visualization of Global MHWs Forecast
In this section, we visualize the global MHWs forecast result of our Ocean-E2E and S2S, as shown in Figure 11-14. We plot the global SSTA field, where darkred parts indicate strong MHWs.
G.4 Visualization of Regional High Resolution Simulation
In this section, we visualize the regional (North Pacific) simulation of $C_{t}-C_{0}$, as shown in Figure 15-18. We plot the regional SSTA field, where darkred parts indicate strong intensifications of the MHWs.
@@ G.4 Visualization of Regional High Resolution Simulation
Figure 10: Global Simulation of MHWs Initialized on October 9, 2020
@@
Figure 14: Global Forecast of MHWs Initialized on October 9, 2020
Figure 15: Regional Simulation of North Pacific Initialized on January 13, 2020
Figure 16: Regional Simulation of North Pacific Initialized on May 12, 2020
Figure 17: Regional Simulation of North Pacific Initialized on July 11, 2020
Figure 18: Regional Simulation of North Pacific Initialized on October 9, 2020

### v3

@@ Conclusions
$\begin{split}C_{t}&=\int_{0}^{t}\left[-\left({\bf u}_{g}(t)+\varepsilon_{gm}{\bf u}_{\theta}({\bf u}_{g}(t),C)\right)\nabla C\right]\mathbf{M}\\ &\quad+\dot{\rm S}_{\theta}({\bf A}(t),C)\,dt+C_{0}\end{split}$ (31)
Time Integration. In this study, we employ the forward Euler method to approximate the time integration of the system, as described by the following discrete updates:
$C_{(i+1)\Delta t}$ $=C_{i\Delta t}+\Big{[}-\big{(}\mathbf{u}_{i\Delta t}+\mathbf{u}_{\theta}(\mathbf{u}_{i\Delta t},C_{i\Delta t})\big{)}\nabla C_{i\Delta t}$
$\quad+\dot{\rm S}_{\theta}(\mathbf{A}_{i\Delta t},\mathbf{A}_{(i+1)\Delta t},C_{i\Delta t})\Big{]}\Delta t,$ (32)
$\mathbf{u}_{(i+1)\Delta t}$ $=\mathcal{M}_{\theta}(\mathbf{u}_{i\Delta t}),$ (33)
$\mathbf{A}_{(i+1)\Delta t}$ $=\mathcal{N}_{\theta}(\mathbf{A}_{i\Delta t}).$ (34)
For global low-resolution simulations at $1/2^{\circ}$ grid spacing, we set the time step $\Delta t=24$ hours and integrate over a total period of $t=4$ days, resulting in 4 discrete steps. This coarser temporal discretization is sufficient given the relatively slower dynamics captured at this scale.
In contrast, for regional high-resolution simulations at $1/12^{\circ}$ grid spacing, the rapid evolution of oceanic processes—such as mesoscale eddies and submesoscale fronts—necessitates a finer time step to maintain numerical stability and accuracy. Accordingly, we adopt $\Delta t=1800$ seconds ($0.5$ hours) and integrate over the same total period of $t=4$ days, amounting to 192 small steps. However, to balance computational efficiency with fidelity, the neural network corrections $\mathbf{u}_{\theta}$ and $\dot{\rm S}_{\theta}$ are applied only every 48 small steps (corresponding to one day).
This hybrid approach is formalized as follows: for each daily interval indexed by $k=0,1,2,3$ (spanning the 4-day period), we perform 48 sub-steps indexed by $j=0,\dots,47$, with the full updates applied at the end of each daily block. For $j=0$ to $47$:
$\begin{split}C_{(k\cdot 48+j+1)\Delta t}&=C_{(k\cdot 48+j)\Delta t}\\ &\quad+\Big{[}-\mathbf{u}_{(k\cdot 48+j)\Delta t}\nabla C_{(k\cdot 48+j)\Delta t}\Big{]}\Delta t.\end{split}$ (35)
Then, at the end of the daily block:
$\mathbf{u}_{(k\cdot 48+48)\Delta t}$ $\leftarrow\mathbf{u}_{(k\cdot 48+48)\Delta t}$ (36)
$+\mathbf{u}_{\theta}(\mathbf{u}_{(k\cdot 48+48)\Delta t},C_{(k\cdot 48+48)\Delta t})$
$C_{(k\cdot 48+48)\Delta t}$ $\leftarrow C_{(k\cdot 48+48)\Delta t}+\dot{\rm S}_{\theta}$ (37)
This design ensures that the computationally intensive neural network evaluations are invoked daily, while the advection term is resolved at high temporal frequency to capture fast-changing dynamics.
Furthermore, as stated in the main text, our lightweight dynamical kernel imposes no substantial computational overhead on the simulations. As demonstrated in Table 8, in regional high-resolution simulations, incorporating the dynamical core results in only a 14% increase in inference time for the Ocean-E2E model.
Table 8: Inference time for Ocean-E2E 30-day forecast (batch size=1, on a single NVIDIA RTX 3090 GPU).
Model Variant Inference Time (ms)
Ocean-E2E w/ Dynamical Kernel 195.29
Ocean-E2E w/o Dynamical Kernel 171.58
Appendix D Appendix D. Details of Our Neural Networks
Multi-scale Convolution
Inspired by (Wu et al. 2024b), this section details the architectures of the neural network modules $\mathcal{M}_{\theta}$ (for modelling ocean surface geostrophic velocity). As shown in Figure 6, both modules leverage the GABlock (Group Attention Block) as their core building unit, designed to efficiently capture multi-scale spatio-temporal features.
@@
The module for approximating the geostrophic velocity, $\mathbf{u}_{g}$, adopts a U-Net-like encoder-decoder architecture. It takes the initial conditions of geostrophic currents $\mathbf{u}_{g}\in\mathbb{R}^{B\times 2\times H\times W}$ as input. The encoder path, composed of $L_{e}$ stages, utilizes downsampling convolutions (e.g., ConvSC) and GABlocks to extract multi-scale features, denoted as $\mathcal{H}_{l}$ for the $l$-th stage. Deeper encoder layers, particularly the bottleneck which processes $\mathcal{H}_{L_{e}}$ with stacked GABlocks, are crucial for capturing non-local spatial dependencies inherent in subgrid effects. The decoder path, with $L_{d}$ stages, symmetrically reconstructs spatial resolution using upsampling convolutions and GABlocks, integrating features from corresponding encoder stages $\mathcal{H}_{L_{e}-l+1}$ via skip connections. A final convolutional layer then maps the decoder’s output to the future geostrophic velocity field $\mathbf{u}_{g}\in\mathbb{R}^{B\times 2\times H\times W}$.
Neural Data Assimilation Network
Inspired by  (Liu et al. 2024), we apply an edge-reparameterization and attention-guided network to project the observation to analysis data, which achieves efficent assimilation for real-time application. It primarily consists of four parts, i.e., a convolutional layer (ConvL), a Kirsch-guided reparameterization module (KRM), a channel attention module (CAM), and a spatial attention module (SAM).
Convolutional Layer
@@
$\begin{aligned} \operatorname{CAM}&=\sigma(\mathrm{MLP}(\mathrm{Avg}(x_{in}^{c}))+\mathrm{MLP}(\mathrm{Max}(x_{in}^{c})))\\ \end{aligned},$ (42)
@@
where $x_{in}^{s}$ is the input of $\operatorname{SAM}$, $\left[~;~\right]$ is exploited to concatenate two types of pooled features.
Kirsch-guided Reparameterization Module
To facilitate efficient assimilation in real‐world MHWs forecasting, we employ the Kirsch‐guided reparameterization module (KRM) with shared parameters, as proposed by Liu et al. (Liu et al. 2024). The KRM is composed of three principal elements: (1) a standard convolutional layer that captures local spatial features; (2) an expand–squeeze convolutional layer that adaptively adjusts the receptive field; and (3) a family of eight directional edge‐detection operators that guide the learning and inference of the convolutional kernels. Specifically, the standard convolutional and expand–squeeze operations can be formulated as follows:
@@
where $\operatorname{perm}(\cdot)$ denotes the permute operation which exchanges the 11st and 22nd dimensions of a tensor, $\operatorname{rep}(\cdot)$ denotes the spatial broadcasting operation. And the predefined eight-direction Kirsch edge filters $K_{i}$ are incorporated into the reparameterization module. To memorize the edge features, the input feature $x_{in}^{k}$ will first be processed by $C\times C\times 1\times 1$ convolution and then use a custom Kirsch filter to extract the feature map gradients in eight different directions. Therefore, the edge information in eight directions can be expressed as follows:
@@
Appendix E Appendix E. Experiments Details
Evaluation Metric
Definition of MHWs. The (Hobday et al. 2016) definition of MHW categories is applied to detect MHWs from original SST field. More precisely, any interval in which the SST anomalies surpass the 90th percentile for a minimum of five days in a row is classified as an MHW event.
@@
${\rm RMSE}(t)=\sqrt{\frac{1}{H\times W}\sum_{i,j}\omega_{i,j}(C_{i,j}(t)-\tilde{C}_{i,j}(t))^{2}}$ (65)
Where $C_{i,j}(t)$ and $\tilde{C}_{i,j}(t)$ is the ground truth and prediction of SST anomaly at lead time $t$, $\omega_{i,j}$ is the weighting coefficient and the subscript $(i,j)$ indicates the data at grid point $(i,j)$ while $H$ and $W$ follow the same definition in Section 2. CSI is the metric that evaluate the forecast ability of extreme events and can be expressed as:
@@
${\rm BIAS}(t)=\frac{1}{H\times W}\sum_{i,j}\omega_{i,j}(C_{i,j}(t)-\tilde{C}_{i,j}(t))$ (67)
Model Training
@@
Appendix F Appendix F. Related Work and Limitations
Numerical Model based Ocean Forecasting
@@
Traditional MHWs forecasting predominantly relies on numerical models solving oceanic primitive equations. These approaches are typically categorized into two types: (1) seasonal forecasting, which predicts large-scale MHW characteristics including onset timing, duration, and intensity evolution (Jacox et al. 2022; Brodie et al. 2023), and (2) sub-seasonal forecasting, which resolves finer spatiotemporal variations at the cost of reduced prediction horizons (Benthuysen et al. 2021; Yu et al. 2024). While physics-based models demonstrate reasonable forecast skill, they face two principal limitations: prohibitive computational costs scaling with model resolution (Xiong et al. 2023; Wang et al. 2024), and systematic underestimation of extreme MHW magnitudes (Jacox et al. 2022).
Deep Learning based Ocean Forecasting
@@
Machine learning (ML) techniques have emerged as promising alternatives for oceanic modeling, offering orders-of-magnitude speedup over conventional methods (Liu and Ma 2024; Hao et al. 2025). Recent efforts have developed spatio-temporal architectures including NMO (Wu et al. 2024a), neXtSIM (Finn et al. 2024), and OceanVP (Shi, Zheng, and Dong 2024) for regional ocean modeling, alongside specialized models for specific oceanic phenomena such as ENSO (Ham, Kim, and Luo 2019; Chen et al. 2025), MJO (Kim et al. 2021; Delaunay and Christensen 2022; Shin et al. 2024), and MHWs (Shi et al. 2024; Jacox et al. 2022; Shu et al. 2025). Building on advances in atmospheric foundation models (Pathak et al. 2022; Bi et al. 2023), the community has recently proposed global ocean foundation models like AI-GOMS (Xiong et al. 2023), Xihe (Wang et al. 2024), and WenHai (Cui et al. 2025).
Deep Learning based Data Assimilation
@@
Deep learning (DL) has shown transformative potential in weather and climate data assimilation (DA), with progress evolving along three directions: (1) Conceptual validation using idealized systems like Lorenz models and shallow water equations (Brajard et al. 2020; Arcucci et al. 2021; Fablet et al. 2021; Legler and Janjić 2022); (2) Component replacement within conventional DA frameworks, including physical parameterization (Hatfield et al. 2021), observation operators (Liang et al. 2022; Stegmann et al. 2022), and optimization cost functions (Melinc and Zaplotnik 2024; Xiao et al. 2024); (3) End-to-end DA system replacement, exemplified by FengWu-Adas and DiffDA that implement complete DL-based assimilation pipelines using ERA5 reanalysis data (Chen et al. 2023a; Huang et al. 2024).
Appendix G Appendix G. Additional Results
Additional Comparison with S2S System
@@
In this section, we are going to show additional comparison results with S2S system. To be specific, we take the S2S analysis field as the groundtruth, and use it to initialize our Ocean-E2E. The atmospheric boundary condition is acquired through one of the state-of-the-art AI-driven weather forecast model: OneForecast (Gao et al. 2025). As shown in Table 9, our model still performs better than S2S.
Visualization of Global MHWs Simulation
In this section, we visualize the global MHWs forecast result of our Ocean-E2E, as shown in Figure 7-9. We plot the global SSTA field, where darkred parts indicate strong MHWs.
Visualization of Global MHWs Forecast
In this section, we visualize the global MHWs forecast result of our Ocean-E2E, as shown in Figure 10-12. We plot the global SSTA field, where darkred parts indicate strong MHWs.
Visualization of Regional High Resolution Simulation
In this section, we visualize the regional (western Atlantic) simulation, as shown in Figure 13-15. We plot the regional SSTA field, where yellow parts indicate MHWs.
@@
Figure 13: Regional Simulation of West Atlantic Initialized on January 13, 2020
Figure 14: Regional Simulation of West Atlantic Initialized on March 13, 2020
Figure 15: Regional Simulation of West Atlantic Initialized on May 12, 2020